OpenStudy (anonymous):

Solve on the interval (0,2pi) 3Sec x-2 = 1 A. pi/6, 5pi/6 B. 2pi/3, 4pi/3 C. pi/3, 5pi/3 D. 0

OpenStudy (anonymous):

rewrite it as 3*1/cosx -2=1 3-2cosx/cosx=1 3-2cosx=cosx 3-3cosx=0 3cosx=3 cosx=1 x= 0